OpenStudy (anonymous):

Order from least to greatest: -2, 4, 2/3, -6, 8/3

OpenStudy (anonymous):

@stamp

OpenStudy (stamp):

What is confusing you?

OpenStudy (anonymous):

the fractions..i understand the numbers and can decifer whcih ones are from least t greatest...but i cant figure out where my fractions are supposed to go

OpenStudy (stamp):

Negative numbers are certainly less than positive numbers. Order all the negative numbers on one side and all the positives on the greater side, then we can start narrowing it down further.

OpenStudy (anonymous):

ok i figured it out..my anser is a

OpenStudy (stamp):

-2, 4, 2/3, -6, 8/3 Least to greatest: -6, -2, 2/3, 8/3, 4

OpenStudy (stamp):

-6 is further away from 0 than -2, so -6 is lesser value (more negative) than -2 is. -6, -2, Now that our negative numbers are handled, we can evaluate 2/3, 8/3, and 4. 2/3 is less than 1 while 8/3 is greater than 1. -6, -2, 2/3, 8/3 Where does the 4 fit in? Well 4 = 4/1 = 12/3 So 12/3 is the greatest of the numbers -6, -2, 2/3, 8/3, 4

OpenStudy (anonymous):

-6 -2 0.6 2.66 4

OpenStudy (anonymous):

2/3=0.6 8/3 =2.66